# SimMerge: Learning to Select Merge Operators from Similarity Signals

### v1

@@ Abstract
Model merging enables multiple large language models (LLMs) to be combined into a single model while preserving performance. This makes it a valuable tool in LLM development, offering a competitive alternative to multi-task training. However, merging can be difficult at scale, as successful merging requires choosing the right merge operator, selecting the right models, and merging them in the right order. This often leads researchers to run expensive merge-and-evaluate searches to select the best merge. In this work, we provide an alternative by introducing SimMerge, a predictive merge-selection method that selects the best merge using inexpensive, task-agnostic similarity signals between models. From a small set of unlabeled probes, we compute functional and structural features and use them to predict the performance of a given 2-way merge. Using these predictions, SimMerge selects the best merge operator, the subset of models to merge, and the merge order, eliminating the expensive merge-and-evaluate loop. We demonstrate that we surpass standard merge-operator performance on 2-way merges of 7B-parameter LLMs, and that SimMerge generalizes to multi-way merges and  [^1] B-parameter LLM merges without retraining. Additionally, we present a bandit variant that supports adding new tasks, models, and operators on the fly. Our results suggest that learning how to merge is a practical route to scalable model composition when checkpoint catalogs are large and evaluation budgets are tight.
@@ 1 Introduction
Model merging is a practical approach for composing checkpoints trained on complementary subsets or domain-specific objectives, producing a single model that aggregates their capabilities with modest additional data or compute. Early approaches such as simple parameter averaging and related weight-space operations improved accuracy and robustness, motivating a broad line of methods for composing models in parameter space [Wortsman et al., 2022; Ilharco et al., 2023; Matena & Raffel, 2022]. More recent work proposes structured or interference-aware rules that make merging more reliable across tasks and initializations [Yadav et al., 2023; Huang et al., 2024; Stoica et al., 2024], as well as perspectives that explain or facilitate merging through subspace matching and permutation alignment [Tam et al., 2023; Ainsworth et al., 2022]. Tooling further standardizes implementations and recipes, making merging accessible at scale [Goddard et al., 2024]. As modern LLM development produces thousands of fine-tuned checkpoints across tasks, modalities, and languages, model composition is becoming a central requirement for scalable AI systems. In this work, we study merging post-trained checkpoints: fine-tuned LLMs derived from a shared pretrained base.
@@
At this scale, the bottleneck shifts from how to merge to how to choose a merge configuration. Given a large catalog, practitioners must decide which models to combine, which operator to apply, and, for multiway merging, in what order. The standard workflow relies on empirical search, where each candidate configuration requires a merge followed by downstream evaluation. Even with efficient operator implementations or data-free heuristics, the cost grows quickly with the number of candidate models, operators, and merge orders, making exhaustive search impractical. Beyond cost, the choice is also sensitive: the same fixed operator can be strong in one domain and harmful in another, so a single global recipe is unreliable in practice.
@@
We address this bottleneck with predictive merge selection. We introduce SimMerge, a lightweight trained selector that predicts the merge operator and, for multiway merges, the merge order using inexpensive, task-agnostic similarity measurements between source checkpoints before merging. Unlike prior work that proposes merge operators or tunes merging hyperparameters [Matena & Raffel, 2022; Yadav et al., 2023; Tam et al., 2023; Ainsworth et al., 2022; Stoica et al., 2024; Huang et al., 2024; Akiba et al., 2024], SimMerge learns the selection decision from pre-merge similarity signals. Using a small set of unlabeled probes, we compute functional and structural features such as KL divergence between model logits, cosine similarity of weights and attention patterns, and $\ell_{2}$ distance in parameter space. These signals are far cheaper than running many merge-and-evaluate trials yet are informative about which configuration will succeed for a given set of models.
@@
SimMerge is not a new merge operator; it learns when to apply existing ones. In the pairwise setting, it selects among linear interpolation [Wortsman et al., 2022], spherical linear interpolation (SLERP) [Shoemake, 1985], and TIES merging [Yadav et al., 2023]. New operators can be added as candidates without changing the framework. The selector is trained offline on previously observed merges: for each candidate pair and operator, we record downstream performance and learn to predict the configuration that maximizes it. At deployment time, SimMerge requires only the checkpoints and a small unlabeled probe set. It performs no downstream evaluation and applies no gradient-based fine-tuning.
@@
Predictive selection is especially valuable for multiway merging, where operator choice alone is insufficient. Common merge operators are not associative, so different merge orders can produce different models. The number of possible orders grows rapidly with the number of models, making merge-and-evaluate search infeasible at scale. SimMerge instead selects a merge plan directly from pre-merge similarity signals. Although trained only on pairwise merges, the similarity-based decision rule extends to multiway settings by scoring plans from the same features, enabling effective three- and four-way merges without extra supervision.
Finally, we consider deployment settings where the model pool and task distribution evolve over time. Offline selectors can drift when new checkpoints or operators appear. We develop an online variant based on a contextual bandit that updates under partial feedback, preserving the low-cost selection step at inference time. Empirically, SimMerge is effective on two- to four-way merges of 7B models across multilingual, code, math, and RAG, and transfers to a  [^1] B model without retraining, reducing expert degradation at scale.
@@
  • We extend pairwise operator selection to multiway merging by scoring entire merge plans using the same pairwise similarity features, representing each plan as an ordered sequence of pairwise merges. We also show this approach scales from 7B to  [^1] B-parameter models without retraining.
  • We introduce an online contextual bandit variant of SimMerge that adapts under partial feedback and supports adding new tasks, models, and operators on-the-fly, allowing for continual deployment in evolving checkpoint catalogs.
  • We conduct extensive empirical evaluations across 2-, 3-, and 4-way merges on code, math, multilingual, RAG, and instruction tasks, demonstrating that SimMerge consistently outperforms the best fixed merge operator. Averaged over the tasks, SimMerge closes $65.0\%$ of the expert-off-domain performance gap, compared to $41.8\%$ for the best single fixed operator.
@@ 2 Methodology
We consider a catalog of domain-specialized language model checkpoints that we would like to combine into a single model that performs well across a set of target tasks. Let $\mathcal{T}$ denote the set of evaluation tasks spanning these domains. The goal is to construct a composite model $\tilde{m}$ that achieves high performance on all tasks in $\mathcal{T}$.
Let $\theta(m)\in\mathbb{R}^{d}$ denote the parameters of model $m$. We work with layerwise binary merge operators $M_{o}$ indexed by $o\in\mathcal{O}$. Given two models $m_{a},m_{b}$ and a mixing coefficient $\alpha\in[0,1]$, an operator produces merged parameters
$\tilde{\theta}=M_{o}\bigl(\theta(m_{a}),\theta(m_{b});\alpha\bigr)\in\mathbb{R}^{d}.$
@@
In all main experiments we use equal-weight merges with $\alpha=0.5$ and take $\mathcal{O}=\{\textsc{Linear}{},\textsc{Slerp}{},\textsc{Ties}{}\}$, where Linear is standard linear interpolation, Slerp is spherical linear interpolation, and Ties (trim, elect sign and merge) is a sign-consistent merge. Formal definitions of operators are detailed in Appendix A.
Given models $\{m_{i_{1}},\dots,m_{i_{k}}\}$, a $k$-way merge plan is an ordered sequence,
$\pi=(m_{i_{1}}\rightarrow m_{i_{2}}\rightarrow\dots\rightarrow m_{i_{k}}),$
realized as k1k-1 sequential applications of binary merge operators. Concretely, for an operator sequence $\mathbf{o}=(o_{1},\dots,o_{k-1})$ with $o_{t}\in\mathcal{O}$, the plan is executed by repeatedly merging the current model with the next model in $\pi$ using $M_{o_{t}}$. Because merge operators are not associative, both the order $\pi$ and the per-step operator choices $\mathbf{o}$ affect the final merged model and its utility on tasks in $\mathcal{T}$. As a result, naive scheme that enumerates all orders and operator choices is quickly intractable: there are $k!$ possible orders and $|\mathcal{O}|^{k-1}$ operator sequences, for a total of $k!\,|\mathcal{O}|^{k-1}$ candidate configurations.
SimMerge is designed to avoid this cost. It precomputes similarity features for each ordered pair of models and task, learns an operator selector from labeled 2-way merges, and then scores and executes multiway merge plans using only these pre-merge signals. Figure 1 gives an end-to-end view of SimMerge. Given a set of checkpoints and small unlabeled probe sets, SimMerge predicts which operator to use at each binary merge step and which merge order to follow (for $k>2$). When we require a single merged model for a task set $\mathcal{T}$, we score candidate configurations by aggregating predicted utility across tasks, using macro-averaging over $t\in\mathcal{T}$.
2.1 Setup and Similarity Features
For every ordered pair of models $(m_{a},m_{b})$ and task $t\in\mathcal{T}$, SimMerge computes a similarity feature vector
$x(m_{a},m_{b},t)\in\mathbb{R}^{m}$
@@ 2.1 Setup and Similarity Features
from an unlabeled probe set and the model weights. These signals are inexpensive to obtain and capture both functional behavior and structural alignment. For each task $t$ we draw an unlabeled probe set $\mathcal{P}_{t}$ from the input distribution of $t$ and compute all data-based metrics on $\mathcal{P}_{t}$. Probe sizes are reported in Section 3.5.
@@
Given an ordered pair $(m_{a},m_{b})$, we first compute measures of functional proximity on $\mathcal{P}_{t}$, including the Kullback-Leibler divergence $D_{\mathrm{KL}}(p_{a}\,\|\,p_{b})$ between predictive distributions, optionally computed per layer and then aggregated, and cosine similarities between intermediate activations $h_{a}^{(\ell)}$ and $h_{b}^{(\ell)}$ at each layer $\ell$, averaged over inputs and layers. We then compute weight-based measures that compare models directly in parameter space, including cosine similarity between flattened parameter vectors $\theta(m_{a})$ and $\theta(m_{b})$, Euclidean distance $\|\theta(m_{a})-\theta(m_{b})\|_{2}$, and the individual norms $\|\theta(m_{a})\|_{2}$ and $\|\theta(m_{b})\|_{2}$. We additionally include cosine similarity of attention patterns as a separate feature channel.
@@
Each metric yields either a scalar or a short sequence over layers or modules. To obtain a fixed-dimensional representation, sequence-valued metrics are summarized using robust statistics such as the mean, median, and selected quantiles, and all summaries are concatenated into $x(m_{a},m_{b},t)$. By default, we append a task encoding $c(t)\in\mathbb{R}^{d_{c}}$ and use
@@
as the input to all learned components. We also evaluate a task-agnostic variant that omits $c(t)$ and report the comparison in Appendix F.1. The improvement from task encoding is modest but consistent, so we use it by default. Full details of the similarity metrics and aggregations are given in Appendix B.
Computation cost. Probe-based features scale linearly with the number of probe tokens and require only forward passes. In our experiments, we use only 10000 probe tokens per task, and we cache per-checkpoint outputs so pairwise comparisons reduce to inexpensive post-processing. Weight-based features require a single pass over the parameters (linear in model size), and the overall cost is far lower than running merge-and-evaluate searches.
@@ 2.2 Predictive Merge Selection
SimMerge replaces merge-and-evaluate search with a prediction policy that uses only pre-merge similarity signals. For a target task $t$, the policy selects a merge operator for each binary merge step. When $k>2$, it also selects the merge order in which those steps are applied.
Pairwise operator selection. For each ordered pair $(m_{a},m_{b})$ of models, each task $t\in\mathcal{T}$, and each operator $o\in\mathcal{O}$, we construct a 2-way merge example as follows. We compute the feature vector $\tilde{x}(m_{a},m_{b},t)$, merge the models with $M_{o}$ at $\alpha=0.5$ to obtain $\tilde{m}_{o}$, and evaluate $\tilde{m}_{o}$ on a held-out validation set for task $t$, yielding a utility score
$s(o;m_{a},m_{b},t):=\mathrm{score}(\tilde{m}_{o},t).$
The pairwise label is defined as
$y^{\star}(m_{a},m_{b},t)=\arg\max_{o\in\mathcal{O}}s(o;m_{a},m_{b},t).$
We also record the per-operator utilities $\big(s(o;m_{a},m_{b},t)\big)_{o\in\mathcal{O}}$.
The pairwise operator selector is a function $f_{\mathrm{op}}:\mathbb{R}^{m+d_{c}}\rightarrow\mathcal{O}$ that takes $\tilde{x}(m_{a},m_{b},t)$ as input and predicts the merge operator that maximizes downstream utility for the pair $(m_{a},m_{b})$ on task $t$. In our offline experiments $f_{\mathrm{op}}$ is instantiated as a lightweight multi-class predictor that outputs a categorical distribution over $\mathcal{O}$ and is trained with a cross-entropy loss to predict $y^{\star}$. Architectural details are given in Section 3.5.
At inference time, given a new pair $(m_{a},m_{b})$ and a task $t$, SimMerge computes $\tilde{x}(m_{a},m_{b},t)$, applies $f_{\mathrm{op}}$ to obtain scores over operators, selects $\hat{o}\in\mathcal{O}$ by taking $\arg\max$, and returns the merged model $\tilde{m}_{\hat{o}}=M_{\hat{o}}(m_{a},m_{b};0.5)$. This replaces an exhaustive loop over all operators with a single guided merge.
Multi-way plan representation and scoring. To handle $k$-way merges, SimMerge scores candidate merge plans using only pairwise similarity features. Given a plan $\pi=(m_{i_{1}}\rightarrow m_{i_{2}}\rightarrow\dots\rightarrow m_{i_{k}})$ and a task $t$, we build an order-aware representation by concatenating feature blocks for each step,
$X(\pi,t)=\big[x(m_{i_{1}},m_{i_{2}},t),\,x(m_{i_{2}},m_{i_{3}},t),\,\dots,\,x(m_{i_{k-1}},m_{i_{k}},t)\big]\oplus c(t)\in\mathbb{R}^{D}.$
This encoding preserves both directionality and step position; different orders induce different $X(\pi,t)$. For steps that involve an intermediate merged model, we approximate the required similarity features by propagating metrics from the original pairwise table as described in Appendix D. All blocks are produced by the same similarity encoder used in the pairwise setting.
A plan scorer $f_{\mathrm{plan}}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ then assigns a scalar score to each candidate plan. In practice $f_{\mathrm{plan}}$ is an MLP that shares the similarity encoder with $f_{\mathrm{op}}$ but has its own final layers. Training $f_{\mathrm{plan}}$ does not require manual annotations. We obtain targets by executing sampled $k$-way plans and evaluating the resulting merged model on task $t$, yielding a utility $U(\pi,t)$, and train $f_{\mathrm{plan}}(X(\pi,t))$ to regress toward $U(\pi,t)$ over a collection of plans and tasks. A theoretical justification for why pairwise-derived supervision can extend to multiway planning is given in Appendix C.
In the multilingual-code-math-RAG setting, for example, this procedure can yield a plan that first merges $m_{\text{code}}$ and $m_{\text{math}}$ using Ties, then adds $m_{\text{multi}}$ using Slerp, and finally incorporates $m_{\text{rag}}$ using Linear, with all choices inferred from pre-merge similarity features rather than by evaluating every candidate plan.
Propagating metrics to intermediate merges. Multi-way plans involve intermediate merged models, such as the result of merging $m_{i_{1}}$ and $m_{i_{2}}$ before comparing to $m_{i_{3}}$. Recomputing similarity signals for every hypothetical intermediate model would largely negate the computational savings of SimMerge. Instead, we approximate similarities involving intermediate merges by propagating metrics from the original pairwise table using convexity properties and standard inequalities. For distributional metrics on probe-set predictions, we treat the intermediate model as an equal-weight mixture of its two parents on the probe set, which yields simple upper bounds for divergences. We fix $\alpha=0.5$ in all experiments to isolate the effects of merge operator and merge order, because tuning $\alpha$ adds a separate hyperparameter search that can confound attribution and reintroduce the expensive search loop.
For predictive distributions $P_{a}$, $P_{b}$, and $Q_{c}$ on the probe set, convexity of $f$-divergences yields upper bounds on terms such as
$\mathrm{KL}\big((1-\alpha)P_{a}+\alpha P_{b}\,\|\,Q_{c}\big)\quad\text{and}\quad\mathrm{KL}\big(Q_{c}\,\|\,(1-\alpha)Q_{a}+\alpha Q_{b}\big)$
in terms of $\mathrm{KL}(P_{a}\,\|\,Q_{c})$ and $\mathrm{KL}(P_{b}\,\|\,Q_{c})$. Similarly, if $\theta_{a},\theta_{b},\theta_{c}$ denote flattened parameter vectors, the triangle inequality gives
$\big\|(1-\alpha)\theta_{a}+\alpha\theta_{b}-\theta_{c}\big\|_{2}\;\leq\;(1-\alpha)\|\theta_{a}-\theta_{c}\|_{2}+\alpha\|\theta_{b}-\theta_{c}\|_{2}.$
Cosine similarities between mixtures and third models are approximated using convex combinations of pairwise cosines and norms, together with bounds derived from Cauchy-Schwarz. These constructions can be applied layerwise or modulewise and the resulting propagated metrics are aggregated with the same statistics used in Section 2.1 and inserted into $X(\pi,t)$ whenever a similarity involving an intermediate merge is required. Full formulas and derivations are provided in Appendix D.
At test time, given a subset of models and a task $t$, SimMerge enumerates or samples a small set of candidate plans, builds $X(\pi,t)$ for each using the original and propagated metrics, scores them with $f_{\mathrm{plan}}$, selects the highest-scoring plan $\hat{\pi}$, and executes $\hat{\pi}$ as a sequence of pairwise merges, choosing the operator at each step via $f_{\mathrm{op}}$.

### v2

@@ Abstract
Model merging combines multiple models into a single model with aggregated capabilities, making it a powerful tool for large language model (LLM) development. However, scaling model merging is challenging: performance depends on the choice of merge operator, model subset, and merge order, often requiring expensive merge-and-evaluate searches. In this work, we introduce SimMerge, a predictive merge-selection method that identifies high-performing merges using inexpensive, task-agnostic similarity signals between models. Given a small set of unlabeled probes, SimMerge extracts functional and structural features to predict the performance of candidate two-way merges, enabling merge operator, order and model subset selection without iterative evaluation. We show that SimMerge consistently outperforms the best fixed merge operator across 7B-parameter LLMs and generalizes to multi-way merges and  [^1] B-parameter LLMs without retraining. We further introduce a bandit variant that supports adding new tasks and operators online. Our results suggest that learning how to merge enables scalable model composition when checkpoint catalogs are large and evaluation budgets are limited.
@@ 1 Introduction
Model merging is a practical approach for composing checkpoints trained on complementary subsets or domain-specific objectives into a single model that aggregates their capabilities with modest additional data or compute. Early weight-space methods such as parameter averaging and related operations improved accuracy and robustness, motivating a broad line of model composition approaches in parameter space [Wortsman et al., 2022; Ilharco et al., 2023; Matena & Raffel, 2022]. More recent work proposes structured or interference-aware rules that make merging more reliable across tasks and initializations [Yadav et al., 2023; Huang et al., 2024; Stoica et al., 2024], as well as perspectives that explain or facilitate merging through subspace matching and permutation alignment [Tam et al., 2023; Ainsworth et al., 2022]. Tooling further standardizes implementations and recipes, making merging accessible at scale [Goddard et al., 2024]. As modern LLM development produces thousands of fine-tuned checkpoints across tasks, modalities, and languages, model composition is becoming a central requirement for scalable AI systems. In this work, we study merging post-trained checkpoints: fine-tuned LLMs derived from a shared pretrained base.
@@
At this scale, the bottleneck shifts from how to merge to how to choose a merge configuration. From a large catalog, practitioners must select (i) which models to combine, (ii) which operator to apply, and (iii) for multiway merging, in what order. The prevailing approach is empirical search: each candidate requires a merge followed by downstream evaluation, and the cost grows rapidly with the number of models, operators, and merge orders. Moreover, the choice is sensitive where an operator that helps in one domain can harm in another, so a single global recipe is unreliable.
@@
We address this bottleneck with predictive merge selection. We introduce SimMerge, a lightweight trained selector that predicts the merge operator and, for multiway merges, the merge order using inexpensive pre-merge similarity signals between source checkpoints. Unlike prior work that proposes merge operators or tunes merging hyperparameters [Matena & Raffel, 2022; Yadav et al., 2023; Tam et al., 2023; Ainsworth et al., 2022; Stoica et al., 2024; Huang et al., 2024; Akiba et al., 2025], SimMerge learns the selection decision from these signals. Using a small set of unlabeled probes, we compute functional and structural features such as KL divergence between model logits, cosine similarity of weights and attention patterns, and $\ell_{2}$ distance in parameter space. These signals are far cheaper than running many merge-and-evaluate search yet predictive of downstream outcomes.
@@
SimMerge is not a new merge operator; it learns when to apply existing ones. In the pairwise setting, it selects among linear interpolation [Wortsman et al., 2022], spherical linear interpolation (SLERP) [Shoemake, 1985], and TIES merging [Yadav et al., 2023]. The selector is trained offline on previously observed merges to predict which configuration maximizes downstream performance; at deployment it requires only checkpoints and probes, with no downstream evaluation or gradient-based fine-tuning.
@@
Predictive selection is especially valuable for multiway merging because common merge operators are not associative: different merge orders can yield different models, and the number of possible orders grows rapidly with the number of checkpoints. Although trained only on pairwise merges, SimMerge extends to multiway settings by scoring candidate merge plans, represented as ordered sequences of pairwise merges, using the same similarity features. This enables effective three- and four-way merges without additional supervision.
We consider deployment settings where the model pool and task distribution evolve over time. Offline selectors can drift when new checkpoints or operators appear. We develop an online variant based on a contextual bandit that updates under partial feedback, preserving the low-cost selection step at inference time. Empirically, SimMerge is effective on two- to four-way merges of 7B models across multilingual, code, math, and RAG, and transfers to a  [^1] B model without retraining, reducing expert degradation at scale.
@@
  • We extend pairwise operator selection to multiway merging by scoring merge plans using the same pairwise similarity features, and show this approach scales from 7B to  [^1] B-parameter models without retraining.
  • We introduce an online contextual bandit variant of SimMerge that adapts under partial feedback and supports adding new tasks, models, and operators on-the-fly for evolving checkpoint catalogs.
  • We conduct extensive evaluations across 2-, 3-, and 4-way merges on code, math, multilingual, RAG, and instruction tasks, showing that SimMerge consistently outperforms the best fixed merge operator. Averaged over the tasks, SimMerge closes $65.0\%$ of the expert-off-domain performance gap, compared to $41.8\%$ for the best single fixed operator.
@@ 2 Methodology
We consider a catalog of post-trained checkpoints $\mathcal{M}=\{m_{1},\dots,m_{n}\}$ derived from a shared pretrained base, and a set of tasks $\mathcal{T}=\{t_{1},\dots,t_{|\mathcal{T}|}\}$. Evaluating a model $m\in\mathcal{M}$ on a task $t\in\mathcal{T}$ yields a scalar utility $U(m,t)$ computed from the task’s evaluation metric on held-out data. Our goal is to construct a single composite model $\tilde{m}$ by merging checkpoints from $\mathcal{M}$ so that $U(\tilde{m},t)$ is high for a target task $t$, or high on average across a set of tasks.
2.1 Merge Operators and Plans
@@
Binary merge operators. Pairwise merges are performed using binary merge operators $o\in\mathcal{O}$. Given two checkpoints $(m_{a},m_{b})$ and a mixing coefficient $\alpha\in[0,1]$, the merged checkpoint is denoted by $o(m_{a},m_{b};\alpha)$. In all experiments we use equal-weight merges with $\alpha=0.5$ and the operator set $\mathcal{O}=\{\textsc{Linear}{},\textsc{Slerp}{},\textsc{TIES}{}\}$. Formal operator definitions are given in Appendix A.
Merge plans. A merge plan specifies an ordered sequence of checkpoints and the merge operator used at each binary step. Given an ordered plan $\pi=(m_{i_{1}},m_{i_{2}},\dots,m_{i_{k}})$, the merged model is obtained by sequentially applying binary merge operators:
$\begin{cases}M_{1}=m_{i_{1}},\\ M_{j}=o_{j}(M_{j-1},m_{i_{j}};\alpha),&j=2,\dots,k,\end{cases}$
where $o_{j}\in\mathcal{O}$ and the final merged model is $\tilde{m}=M_{k}$.
Many merge operators are not associative, including Slerp and TIES, so both the order in $\pi$ and the per-step operator choices $\{o_{j}\}$ can affect the resulting model and its utility $U(\tilde{m},t)$. The number of candidate plans grows rapidly with the number of checkpoints: for a fixed subset of $k$ models there are $k!$ possible orders and $|\mathcal{O}|^{k-1}$ operator sequences, making merge-and-evaluate search infeasible at scale. SimMerge addresses this bottleneck by selecting operators and merge plans from pre-merge similarity signals.
2.2 Similarity Features
@@
For each task $t\in\mathcal{T}$ we draw a small unlabeled probe set $\mathcal{P}_{t}$ from the input distribution of $t$. For every ordered pair of checkpoints $(m_{a},m_{b})$ and task $t$, SimMerge constructs a feature vector $x(m_{a},m_{b},t)\in\mathbb{R}^{m}$ using $\mathcal{P}_{t}$ and the checkpoint weights. These metrics are inexpensive to compute and capture both functional behavior and structural alignment. (Probe sizes are reported in Section 3.5.)
@@
We first compute functional similarity signals on $\mathcal{P}_{t}$, including the KL divergence between predictive distributions, $D_{\mathrm{KL}}(p_{a}\,\|\,p_{b})$, and cosine similarity between intermediate activations $h_{a}^{(\ell)}$ and $h_{b}^{(\ell)}$ at each layer $\ell$, averaged over inputs and layers. We then compute weight-based measures that compare checkpoints directly in parameter space, including cosine similarity between flattened parameter vectors, Euclidean distance between parameters, and parameter norms. We additionally include cosine similarity of attention patterns as a separate feature channel.
@@
Many metrics yield either a scalar or a short sequence over layers or modules. To obtain a fixed-dimensional representation, we summarize sequence-valued metrics using mean, median, and selected quantiles, and concatenate all summaries into $x(m_{a},m_{b},t)$. By default, we append a task encoding $c(t)\in\mathbb{R}^{d_{c}}$ and use
@@
as input to all learned components. We also evaluate a task-agnostic variant that omits $c(t)$ and report the comparison in Appendix F.1. The improvement from task encoding is modest but consistent, so we use it by default. Full details of the similarity metrics and aggregation procedures are provided in Appendix B.
Computation cost. Probe-based metrics require only forward passes over small probe sets. We cache per-checkpoint probe outputs so pairwise comparisons reduce to inexpensive post-processing. Weight-based metrics require a single pass over parameters, and the overall cost is far lower than running merge-and-evaluate searches.
@@ 2.3 Predictive Merge Selection
Using the precomputed similarity features, SimMerge predicts merge operators and scores candidate merge plans without executing merges during selection. Throughout this section we fix the mixing coefficient to $\alpha=0.5$ and omit it from the notation, writing $o(m_{a},m_{b})$ for $o(m_{a},m_{b};0.5)$.
2.3.1 Pairwise utility prediction.
SimMerge is built around a learned utility predictor that estimates the performance of a merge configuration for a given task. In the pairwise setting, for an ordered pair of checkpoints $(m_{a},m_{b})$ and task $t\in\mathcal{T}$, we define a utility predictor
$f_{\mathrm{plan}}:\mathbb{R}^{m+d_{c}}\rightarrow\mathbb{R}^{|\mathcal{O}|},$
which predicts the utility of merging $m_{a}$ and $m_{b}$ under each merge operator $o\in\mathcal{O}$. The predictor takes as input pre-merge similarity features $\tilde{x}(m_{a},m_{b},t)$ and is trained on offline two-way merges, where each operator is applied and the downstream utility $U(o(m_{a},m_{b}),t)$ is observed. Training minimizes a regression loss toward the observed utilities for all operators:
$\widehat{U}(o(m_{a},m_{b}),t)=f_{\mathrm{plan}}(\tilde{x}(m_{a},m_{b},t))_{o}.$
At inference time, we select the operator with the highest predicted utility,
$\hat{o}=\arg\max_{o\in\mathcal{O}}\widehat{U}(o(m_{a},m_{b}),t),$
replacing exhaustive merge-and-evaluate search.
2.3.2 Multi-way merge plan scoring.
Multi-way merging generalizes this formulation by selecting an entire merge plan that maximizes predicted utility. A $k$-way merge plan $\pi=(m_{i_{1}},\ldots,m_{i_{k}})$ is represented as an ordered sequence of binary merge steps. SimMerge estimates the utility of a plan by recursively predicting the utility of each intermediate merge, using the same plan scorer to select operators at each step.
Because intermediate merged models are not explicitly constructed during scoring, similarity features involving an intermediate model $M_{1:k-1}$ are approximated using features propagated from the original pairwise similarity table (see Appendix D). Different merge orders induce different sequences of predicted utilities, allowing SimMerge to capture the effects of both operator choice and merge order.
At test time, SimMerge enumerates or samples a small set of candidate merge plans, estimates their predicted utilities, and selects the plan with the highest predicted utility:
$\hat{\pi}=\arg\max_{\pi}\widehat{U}(\pi,t).$
The selected plan is then executed as a sequence of binary merges with the step-wise operators $\{\hat{o}_{j}\}$ chosen by the pairwise predictor.